How many solutions does the system x + y = 3 and 2x + 2y = 5 have?
Answer by mananth(16946)

x + y = 3
y=-x+3..................(1)
and 2x + 2y = 5
2y=-2x+5
y=-x+5/2..................(2)
the slopes of equations (10) & (2) are same and the intercepts are different
so the lines are parallel.
it has no solution